§The Chocolate Emporium in Nottingham makes and sells luxury chocolates.

If the price of a box of chocolates increases from £6.99 to £7.49 and demand falls from 862 units per week to 784 units what is the priced elasticity of demand (P.E.D.) for chocolates


1
Expert's answer
2021-05-05T13:34:14-0400

The price elasticity of demand (P.E.D.) for chocolates is:

"Ed = \frac{784-862} {7.49-6.99} ×\frac{7.49+6.99} {784+862} = -1.37."

So, the demand is elastic.